Two bicycles begin at the same point on a circular path traveling in the same direction. The rate of one cyclist is 5 mph faster than the other. If it takes faster cyclist 3 hours to "lap" the slower cyclist, then determine the distance around the path. [Hint: The difference of the distance traveled equals one lap]. If the slower bicycle travels at 10 mph , how many laps did each bicycle make?
Answer by ankor@dixie-net.com(22740)

Two bicycles begin at the same point on a circular path traveling in the same direction.
The rate of one cyclist is 5 mph faster than the other.
If it takes faster cyclist 3 hours to "lap" the slower cyclist, then determine the distance around the path.
:
The relative speed between the cyclists is 5 mph, and it takes 3 hrs to lap:
3 * 5 = 15 mi is the distance around the circular path
:
If the slower bicycle travels at 10 mph, how many laps did each bicycle make?
:
In 3 hrs the slower bike would travel: 3*10 = 30 mi; 30/15 = 2 laps
In 3 hrs the faster bike would travel: 3*15 = 45 mi; 45/15 = 3 laps